OpenStudy (anonymous):

Need help urgently. (3sqr2 + 1) (2sqr2 -3) sqr represents the radical sign

OpenStudy (anonymous):

[9-7\sqrt{2}\]

OpenStudy (anonymous):

foil problem

OpenStudy (anonymous):

\[(6*2)-9\sqrt{2}+2\sqrt{2}-3\]

OpenStudy (anonymous):

\[12-7\sqrt{2}-3\]

OpenStudy (anonymous):

\[9-7\sqrt{2}\]

OpenStudy (anonymous):

i know it's a foil problem but i'm not getting it. how come you got (6*2) - etc. i'm lost